# Efficient and high-performance routing of lattice-surgery paths on three-dimensional lattice

### v3

@@ 3.5.2 Dijkstra’s algorithm extended in the time direction (3D Dijkstra)
As another method utilizing time direction, we propose 3D Dijkstra. The basic scheme of 3D Dijkstra is the same as that of 3D BFS. The difference is that 3D Dijkstra assigns the weight to the voxels and finds a path that minimizes the sum of weights. We assign large weights as the height of voxels increases. This weighting prioritizes paths with lower voxels over those with higher voxels, and we can expect that the 3D lattice can be used more densely by packing paths from the bottom. To find the shortest path on a weighted lattice, it is efficient to use a variant of Dijkstra’s algorithm for sparse graphs [fredman1987fibonacci]: it runs in $O(V\log V+E)$ time for a general graph with $V$ vertices and $E$ edges.
We heuristically choose the weights as $2^{h}$, where $h$ is the height of the voxel. The reason for using exponential weights is that the ratio of the costs of adjacent layers is constant, regardless of the current height. The reason for adopting $2$ as the base is that the scheduling result was found to be efficient when measured and that the algorithm runs fast because it can be implemented with integer bit shifts.
@@ 3.5.3 Dijkstra projection
As shown later in the numerical evaluation parts, the methods 3D BFS and 3D Dijkstra are high-performance but time-consuming for practical problem instances, while they are polynomial-time algorithms. They are especially time-consuming for deep circuits because their time complexity of path search depends on the height of the 3D lattice $h$, as shown in Tab. 1. This motivates us to propose a more time-efficient algorithm, Dijkstra projection. It searches for a 2D path on a 2D lattice and then obtains a 3D path by stacking the 2D path in the time direction like a projection. While 3D BFS and 3D Dijkstra search for a path in a 3D lattice, Dijkstra projection searches for a path only in a 2D lattice, which is likely to speed up the scheduling process.
Dijkstra projection determines a pre-projection path in a similar way to 3D Dijkstra. First, each cell of the qubit plane is weighted exponentially with respect to the heights of the paths already stacked. On the weighted 2D lattice, Dijkstra’s algorithm searches for the path with the minimum sum of the weights. The path is then stacked in the 3D lattice along the time axis with its 2D shape preserved and adjacent voxels of the path sharing a face. Fig. 10 shows an example of stacking a path.
@@ 3.5.4 Look-ahead Dijkstra projection
Instruction look-ahead, introduced in Sec. 2.4, is a technique employed to efficiently manage the set of currently executable instructions. While this technique was initially utilized to mitigate scheduling delays caused by routing failures, we can also leverage it to provide the proposed algorithms with greater flexibility in selecting the next instruction for execution. Specifically, we integrate it with Dijkstra projection and call it look-ahead Dijkstra projection.
@@
Finally, we describe the modifications required to integrate this optimization into Dijkstra projection. A similar approach applies to our other two algorithms. The integration involves three primary changes:
@@
    Then, we modify Alg. 4 by replacing Line 8 with GetExecutableIndex and inserting UpdateDependencyGraph immediately after Line 15. It is noteworthy that these changes barely affect the total time complexity, since both functions run in $O(\log m)$ time, which is negligible compared to $O(n)$ time required for path searching.
@@ 3.5.5 Modifications to meet the kink condition
We now discuss how we can let algorithms find paths satisfying kink conditions. Here, we focus on Dijkstra projection and show how to adjust a 3D path of lattice surgery so that the parity of kinks becomes even. Our approach is to add a process to modify 3D paths after projection. We propose a procedure that flips the kink parity of a projected path, provided that it possesses at least one 90-degree corner. Note that this condition is satisfied when modifying a projected path for an $XX$ or $ZZ$ measurement that has an odd number of kinks. This procedure is employed in the numerical evaluation because it runs in time linear in the path length, which makes it highly efficient.
The core concept for path modification involves lifting corner voxels, as illustrated in Figs. 11 and 12. However, the challenge lies in the fact that this modification might inadvertently create or erase another kinks, thereby failing to flip the parity. To circumvent this challenge, we propose the following procedure that works on the first corner in the path.
@@ 3.6 Time complexity analysis
This section derives the time complexities shown in Tabs. 1 and 2. We first analyze the time complexity of the path finding subroutines shown in Tab. 1. We begin by recalling that BFS and Dijkstra’s algorithm [fredman1987fibonacci] find a shortest path in unweighted and weighted graph, respectively. Their respective running times for a graph with $V$ vertices and $E$ edges are $O(V+E)$ and $O(V\log V+E)$. Since a 2D lattice with $n$ cells has $V=O(n)$ vertices and $E=O(n)$ edges, the path search subroutines of BFS and Dijkstra projection run in $O(n)$ and $O(n\log n)$ time, respectively. Similarly, since a 3D lattice with height $h$ has $V=O(nh)$ vertices $E=O(nh)$ edges, the path search subroutines of 3D BFS and 3D Dijkstra are $O(nh)$ and $O(nh\log nh)$, respectively.
@@
Regarding BFS, the bottleneck of the iterative procedure is FindShortestPath, which runs in $O(n)$ time. Since the failure in routing resets the qubit plane, the numbers of calling FindShortestPath can be bounded above by $2m$. Thus, the total runtime complexity is $O(nm)$.
We then analyze the runtime of look-ahead BFS, which does not admit a favorable bound. Unlike in BFS, the number of times FindShortestPath is called can only be bounded as $O(nm)$. This bound is obtained because the outer loop iterates $O(m)$ times and the inner loop iterates $\left|L\right|=O(n)$ times. Furthermore, this bound may be tight in a pathological case where a single path partitions the qubit plane and subsequent executable instructions attempt to connect two qubits lying on different components. Thus, by incorporating the runtime of instruction look-ahead, the total runtime complexity can be bounded by $O(n^{2}m+m\log m)$. Despite this poor theoretical bound, the algorithm runs almost as fast as BFS in the numerical experiments.
Next, we analyze 3D BFS and 3D Dijkstra simultaneously as they share the algorithm except for the path search subroutine. Since the height $h$ of the 3D lattice can be evaluated as $O(m)$, the path search subroutine runs in $O(nm)$ for 3D BFS and $O(nm\log nm)$ for 3D Dijkstra. Since the routing subroutine never fails to find a path, the number of iteration is $m$, resulting in the total runtime of $O(nm^{2})$ for 3D BFS and $O(nm^{2}\log nm)$ for 3D Dijkstra.
Lastly, we analyze the time complexity of Dijkstra projection and look-ahead Dijkstra projection. As shown in Alg. 4, the overall time complexity of Dijkstra projection is dominated by the function FindShortestPathFromHeight, which executes in $O(n\log n)$ time. We note that the function LiftPath will not be a bottleneck, since it runs in time linear to the path length $\left|p\right|$, even when modifying the kink parity to satisfy the condition. Given that the path length $\left|p\right|$ can be bounded above by $n$, this procedure runs in $O(n)$ time. Consequently, we have a total time complexity of $O(nm\log n)$ for the original algorithm and $O(nm\log n+m\log m)$ when combined with instruction look-ahead.
@@ 4 Performance evaluation
In this section, we evaluate the proposed methods (3D BFS, 3D Dijkstra, Dijkstra projection, and look-ahead Dijkstra projection), and compare them with the baseline methods (BFS and look-ahead BFS). We used random circuits and circuits compiled from quantum phase estimation algorithms for benchmarking. For quantum phase estimation, we first converted the circuit into an instruction sequence consisting of Pauli-$XX$ and $ZZ$ measurements and compared the performance of our methods in Sec. 4.2. We also compared our methods with the existing algorithm for large-scale compilation, the Edge-Disjoint Paths Compilation (EDPC) algorithm [beverland2022surface], in Sec. 4.3. To satisfy their requirements and reproduce their experimental settings, we converted the circuit into another instruction sequence consisting of CNOT operations and $T$-gates and computed the scheduling results on their qubit plane pattern. See each section and Appendix B.4 for further details of these conversions and settings.
@@ 4.1 Random circuits
For the random circuit benchmark, we set the number of total instructions to 1000, and the two logical qubits connected by each instruction were chosen uniformly at random. The data cells were placed in locations where the rows and columns are both even-numbered, as shown in Fig. 1. The number of logical qubits with data lined up on one side of the square qubit plane was defined as qubit plane size. For example, Fig. 1 corresponds to the example of qubit plane size equal to four. The performance was evaluated by changing the qubit plane size from $2$ to $20$. For evaluating the throughput and the execution time, we evaluated methods with ten random circuits generated with different seed values and averaged the results.
@@
In this figure, each color corresponds to the performance of the proposed and baseline methods. Here, we emphasize that the performance of the four methods shown as dotted lines in the figure, 3D BFS, 3D Dijkstra, Dijkstra projection (without kink condition), and look-ahead Dijkstra projection (without kink condition), do not satisfy the kink conditions. They are shown as a performance reference, and only the plots of baseline methods, Dijkstra projection (with kink condition), and look-ahead Dijkstra projection (with kink condition) are valid. Since the optimization of Dijkstra projection by instruction look-ahead hardly impacted throughput or execution time when evaluated on random circuits, we refer to both the original and optimized versions simply as Dijkstra projection throughout this setting.
The methods 3D Dijkstra and Dijkstra projection had the best throughput, followed in order by 3D BFS, look-ahead BFS, and BFS. The small discrepancy between Dijkstra projection and 3D Dijkstra suggests that the limited search space of Dijkstra does not significantly affect the quality of the obtained solution. We find that the penalty of considering the kink condition is negligible since the performance difference of Dijkstra projection with and without considering kink conditions is small. In terms of the increase in throughput as the qubit plane size increased, the best baseline method, look-ahead BFS, showed a little plateauing, while Dijkstra projection showed no such trend and increased at a similar rate. The throughput of Dijkstra projection was on average $1.5$ times higher than that of look-ahead BFS, but the improvement is likely to be even greater as the qubit plane size increases based on the above observation.
@@
We observed that Dijkstra projection and two baseline methods show better scaling compared to 3D BFS and 3D Dijkstra. This is because the first three methods search on the 2D lattice, while the other two search on the 3D lattice. These results are consistent with the expected properties of the algorithm design. We also observed that there is almost no difference between the execution times of Dijkstra projection with and without considering kink conditions. Thus, the time for adjusting kink parity is negligible compared to the path-finding part.
@@ 4.2 SELECT circuits
We then evaluated the performance for practical applications. Here, we chose SELECT circuit as a benchmark target. The SELECT operation is one of the bottleneck components in quantum phase estimation algorithms using qubitization [low2019hamiltonian, babbush2018encoding, yoshioka2022hunting]. Its action is defined from Hamiltonian $H=\sum_{i=0}^{L-1}\alpha_{i}P_{i}$ as follows:
@@
In the following evaluation, we compare four methods, BFS, look-ahead BFS, Dijkstra projection, and look-ahead Dijkstra projection. We omitted the results of 3D BFS and 3D Dijkstra since they cannot finish the compilation within 1 hour.
@@
Dijkstra projection and its optimized variant show better throughput than the two baseline methods. Compared to look-ahead BFS, look-ahead Dijkstra projection shows improvement by a factor of $1.6$ for SELECT 0 and $3.8$ for SELECT 5. We observed higher throughput improvement for SELECT 5, which is an expected result considering that SELECT 5 has a structure like a parallel circuit and leaves much room for parallelization. We can see that the degradation by considering kink conditions is at most 5%, which is not negligible, but does not lose the advantage of Dijkstra projection.
@@
The order of execution times of the four methods for the SELECT circuits was basically the same as for the random circuit. On the other hand, look-ahead Dijkstra projection demonstrated a significant speedup compared to the original Dijkstra projection when applied to SELECT 5. Further profiling indicated that these execution times were dominated by path searching, which strongly suggests that the instruction look-ahead technique results in substantially earlier termination of the path search subroutine. This is likely because fetching the instruction with the lowest height by instruction look-ahead flattens the current scheduling outcome with less biased height distribution, thereby shortening the subsequent lattice-surgery paths and the time required to find them.
All four methods finished in practical execution time. The additional time for considering kink conditions is at most 2%. Thus the time for considering kink conditions is also negligible in this case. Also, when applied to the SELECT 0 circuit, 3D BFS and 3D Dijkstra cannot finish within an hour while Dijkstra projection finishes within a few seconds, again indicating that Dijkstra projection excels in execution time among the methods that utilize the time direction.
@@ 4.3 Comparison with the existing methods
In the following evaluation, four of our methods—BFS, look-ahead BFS, Dijkstra projection, and look-ahead Dijkstra projection—were compared with the Edge-Disjoint Paths Compilation (EDPC) algorithm [beverland2022surface]. We adopted a problem setting that satisfies the requirements of the EDPC algorithm. Specifically, the instruction sequence consists solely of CNOT gates and $T$-gates. Each $T$-gate is realized via $T$-gate teleportation between the target qubit and a magic state factory aligned at the perimeter of the qubit plane. (For further details, see Appendix B.4.) Under this setting, SELECT 0 required $17$ qubit plane size and 37,85637{,}856 instructions, and SELECT 5 required $30$ qubit plane size and 39,05039{,}050 instructions.
Table 5: Throughput evaluation on SELECT circuits. The grayed rows annotated with an asterisk represent reference values that disregard the kink condition and assume CNOT operations complete via a 2D path in $1$ code beat. Although this assumption makes their outputs infeasible, halving these values will provide the corresponding feasible values. The remaining methods adhere to the kink condition and yield a feasible solution.
@@
The throughput of these methods is shown in Tab. 5. For the performance of the EDPC algorithm, it lies between that of BFS and look-ahead BFS. This is due to their different strategy of managing instruction dependency. Their algorithm initially applies topological sorting to the instruction dependency graph and repeatedly removes the instructions that has no dependent instructions, thereby decomposing the instructions into a sequence of instruction sets consisting of mutually independent instructions. This partitioning enables a parallel compilation of the instruction sets and facilitates designing a scheduler that handles many types of instructions, whereas it also worsens the scheduling performance by effectively adding “barriers” in the circuit. We consider this to be one of the reasons why look-ahead BFS outperforms EDPC, although their routing strategy is more sophisticated than our baseline methods.
Although the overall trends of our methods are consistent with the results in Sec. 4.2, it is worth mentioning that the instruction look-ahead technique proved exceptionally effective in this instance. This almost closed the gap between look-ahead BFS and Dijkstra projection, and resulted in an outstanding performance by look-ahead Dijkstra projection.
This gap reduction can also be attributed to the difference in CNOT gate handling. The former experiment decomposes CNOT gates into $ZZ$ and $XX$ measurements, which partially yields stair-like structures and is likely advantageous for 3D routing. However, we emphasize that the same benefits can also be observed in this experiment when we compare feasible throughputs by halving the grayed ones: Dijkstra projection showed $1.8$-fold improvement over look-ahead BFS, and look-ahead Dijkstra projection demonstrated $3.6$-fold improvement.
@@
The execution times of the methods are shown in Tab. 6. The subtantial execution time of EDPC may stem from several reasons. We first note that their algorithm was implemented in Julia whereas ours are implemented in C++. Although we used @btime to warm up the program and exclude just-in-time compilation time, this difference likely introduced overhead to their execution time. Another factor is their use of maximum flow subroutines in $T$-gate teleportation, which may have incurred high computational cost.

### v4

@@ 3.5.2 Dijkstra’s algorithm extended in the time direction (3D Dijkstra)
As another method utilizing time direction, we propose the 3D Dijkstra. The basic scheme of the 3D Dijkstra is the same as that of the 3D BFS. The difference is that the 3D Dijkstra assigns the weight to the voxels and finds a path that minimizes the sum of weights. We assign large weights as the height of voxels increases. This weighting prioritizes paths with lower voxels over those with higher voxels, and we can expect that the 3D lattice can be used more densely by packing paths from the bottom. To find the shortest path on a weighted lattice, it is efficient to use a variant of Dijkstra’s algorithm for sparse graphs [16]: it runs in $O(V\log V+E)$ time for a general graph with $V$ vertices and $E$ edges.
We heuristically choose the weights as $2^{h}$, where $h$ is the height of the voxel. The reason for using exponential weights is that the ratio of the costs of adjacent layers is constant, regardless of the current height. The reasons for adopting $2$ as the base are twofold: the scheduling result was found to be efficient when measured, and the algorithm runs fast because it can be implemented with integer bit shifts.
@@ 3.5.3 Dijkstra projection
As shown later in the numerical evaluation parts, the 3D BFS and 3D Dijkstra methods are high-performance but time-consuming for practical problem instances, while they are polynomial-time algorithms. They are especially time-consuming for deep circuits because their time complexity of path search depends on the height of the 3D lattice $h$, as shown in Tab. 1. This motivates us to propose a more time-efficient algorithm, the Dijkstra projection. It searches for a 2D path on a 2D lattice and then obtains a 3D path by stacking the 2D path in the time direction, like a projection. While the 3D BFS and the 3D Dijkstra search for a path in a 3D lattice, the Dijkstra projection searches for a path only in a 2D lattice, which is likely to speed up the scheduling process.
The Dijkstra projection determines a pre-projection path in a similar way to the 3D Dijkstra. First, each cell of the qubit plane is weighted exponentially with respect to the heights of the paths already stacked. On the weighted 2D lattice, Dijkstra’s algorithm searches for the path with the minimum sum of the weights. The path is then stacked in the 3D lattice along the time axis, preserving its 2D shape while ensuring that adjacent voxels of the path share a face. Fig. 10 shows an example of stacking a path.
@@ 3.5.4 Look-ahead Dijkstra projection
Instruction look-ahead, introduced in Sec. 2.4, is a technique employed to efficiently manage the set of currently executable instructions. While this technique was initially utilized to mitigate scheduling delays caused by routing failures, we can also leverage it to provide the proposed algorithms with greater flexibility in selecting the next instruction for execution. Specifically, we integrate it with the Dijkstra projection and call it look-ahead Dijkstra projection.
@@
Finally, we describe the modifications required to integrate this optimization into the Dijkstra projection. A similar approach applies to our other two algorithms. The integration involves three primary changes:
@@
    Then, we modify Alg. 4 by replacing Line 8 with GetExecutableIndex and inserting UpdateDependencyGraph immediately after Line 15. It is noteworthy that these changes barely affect the total time complexity, since both functions run in $O(\log m)$ time, which is negligible compared to the $O(n)$ time required for path searching.
@@ 3.5.5 Modifications to meet the kink condition
We now discuss how we can let algorithms find paths satisfying kink conditions. Here, we focus on the Dijkstra projection algorithm and its look-ahead variant, and show how to adjust a 3D path of lattice surgery so that the parity of kinks becomes even. Our approach is to add a process to modify 3D paths after projection. We propose a procedure that flips the kink parity of a projected path, provided that it possesses at least one 90-degree corner. Note that this condition is satisfied when modifying a projected path for an $XX$ or $ZZ$ measurement that has an odd number of kinks. This procedure is employed in the numerical evaluation because it runs in time linear in the path length, which makes it highly efficient.
The core concept for path modification involves lifting corner voxels, as illustrated in Figs. 11 and 12. However, the challenge lies in the fact that this modification might inadvertently create or erase another kink, thereby failing to flip the parity. To circumvent this challenge, we propose the following procedure that works on the first corner in the path.
@@ 3.6 Time complexity analysis
This section derives the time complexities shown in Tabs. 1 and 2. We first analyze the time complexity of the path finding subroutines shown in Tab. 1. We begin by recalling that the BFS algorithm and Dijkstra’s algorithm [16] find a shortest path in an unweighted and weighted graph, respectively. Their respective running times for a graph with $V$ vertices and $E$ edges are $O(V+E)$ and $O(V\log V+E)$. Since a 2D lattice with $n$ cells has $V=O(n)$ vertices and $E=O(n)$ edges, the path search subroutines of the BFS and the Dijkstra projection run in $O(n)$ and $O(n\log n)$ time, respectively. Similarly, since a 3D lattice with height $h$ has $V=O(nh)$ vertices and $E=O(nh)$ edges, the path search subroutines of the 3D BFS and the 3D Dijkstra are $O(nh)$ and $O(nh\log nh)$, respectively.
@@
Regarding the BFS, the bottleneck of the iterative procedure is FindShortestPath, which runs in $O(n)$ time. Since the failure in routing resets the qubit plane, the number of calls to FindShortestPath can be bounded above by $2m$. Thus, the total runtime complexity is $O(nm)$.
We then analyze the runtime of the look-ahead BFS, which does not admit a favorable bound. Unlike in the BFS, the number of times FindShortestPath is called can only be bounded as $O(nm)$. This bound is obtained because the outer loop iterates $O(m)$ times and the inner loop iterates $\left|L\right|=O(n)$ times. Furthermore, this bound may be tight in a pathological case where a single path partitions the qubit plane and subsequent executable instructions attempt to connect two logical qubits lying on different components. Thus, by incorporating the runtime of instruction look-ahead, the total runtime complexity can be bounded by $O(n^{2}m+m\log m)$. Despite this poor theoretical bound, the algorithm runs almost as fast as the BFS in the numerical experiments.
Next, we analyze the 3D BFS and the 3D Dijkstra simultaneously as they share the algorithm except for the path search subroutine. Since the height $h$ of the 3D lattice can be evaluated as $O(m)$, the path search subroutine runs in $O(nm)$ for the 3D BFS and $O(nm\log nm)$ for the 3D Dijkstra. Since the routing subroutine never fails to find a path, the number of iterations is $m$, resulting in the total runtime of $O(nm^{2})$ for the 3D BFS and $O(nm^{2}\log nm)$ for the 3D Dijkstra.
Lastly, we analyze the time complexity of the Dijkstra projection and its look-ahead variant. As shown in Alg. 4, the overall time complexity of the Dijkstra projection is dominated by the function FindShortestPathFromHeight, which executes in $O(n\log n)$ time. We note that the function LiftPath will not be a bottleneck, since it runs in time linear to the path length $\left|p\right|$, even when modifying the kink parity to satisfy the condition. Given that the path length $\left|p\right|$ can be bounded above by $n$, this procedure runs in $O(n)$ time. Consequently, we have a total time complexity of $O(nm\log n)$ for the original algorithm and $O(nm\log n+m\log m)$ when combined with instruction look-ahead.
@@ 4 Performance evaluation
In this section, we evaluate the proposed methods (3D BFS, 3D Dijkstra, Dijkstra projection, and look-ahead Dijkstra projection) and compare them with the baseline methods (BFS and look-ahead BFS). We used random circuits and circuits compiled from quantum phase estimation algorithms for benchmarking. For quantum phase estimation, we first converted the circuit into an instruction sequence consisting of Pauli-$XX$ and $ZZ$ measurements and compared the performance of our methods in Sec. 4.2. We also compared our methods with the existing algorithm for large-scale compilation, the Edge-Disjoint Paths Compilation (EDPC) algorithm [4], in Sec. 4.3. To satisfy their requirements and reproduce their experimental settings, we converted the circuit into another instruction sequence consisting of CNOT operations and $T$-gates and computed the scheduling results on their qubit plane pattern. See each section and Appendix B.4 for further details of these conversions and settings.
@@ 4.1 Random circuits
For the random circuit benchmark, we set the number of total instructions to 1000, and the two logical qubits connected by each instruction were chosen uniformly at random. The data cells were placed in locations where the rows and columns are both even-numbered, as shown in Fig. 1. The number of logical qubits with data lined up on one side of the square qubit plane was defined as qubit plane size. For example, Fig. 1 corresponds to the example of a qubit plane size equal to four. The performance was evaluated by changing the qubit plane size from $2$ to $20$. For evaluating the throughput and the execution time, we evaluated methods with ten random circuits generated with different seed values and averaged the results.
@@
In this figure, each color corresponds to the performance of the proposed and baseline methods. Here, we emphasize that the performance of the four methods shown as dotted lines in the figure, 3D BFS, 3D Dijkstra, Dijkstra projection (without kink condition), and look-ahead Dijkstra projection (without kink condition), do not satisfy the kink conditions. They are shown as a performance reference, and only the plots of the baseline methods, the Dijkstra projection (with kink condition), and the look-ahead Dijkstra projection (with kink condition) are valid. Since the optimization of the Dijkstra projection by instruction look-ahead hardly impacted throughput or execution time when evaluated on random circuits, we refer to both the original and optimized versions simply as Dijkstra projection throughout this setting.
The methods 3D Dijkstra and Dijkstra projection had the best throughput, followed in order by 3D BFS, look-ahead BFS, and BFS. The small discrepancy between the Dijkstra projection and the 3D Dijkstra suggests that the limited search space of the Dijkstra projection does not significantly affect the quality of the obtained solution. We find that the penalty of considering the kink condition is negligible since the performance difference of the Dijkstra projection with and without considering kink conditions is small. In terms of the increase in throughput as the qubit plane size increased, the best baseline method, look-ahead BFS, showed a little plateauing, while the Dijkstra projection showed no such trend and increased at a similar rate. The throughput of the Dijkstra projection was on average $1.5$ times higher than that of the look-ahead BFS, but the improvement is likely to be even greater as the qubit plane size increases based on the above observation.
@@
We observed that the Dijkstra projection and two baseline methods show better scaling compared to the 3D BFS and the 3D Dijkstra. This is because the first three methods search on the 2D lattice, while the other two search on the 3D lattice. These results are consistent with the expected properties of the algorithm design. We also observed that there is almost no difference between the execution times of the Dijkstra projection with and without considering kink conditions. Thus, the time for adjusting kink parity is negligible compared to the path-finding part.
@@ 4.2 SELECT circuits
We then evaluated the performance for practical applications. Here, we chose the SELECT circuit as a benchmark target. The SELECT operation is one of the bottleneck components in quantum phase estimation algorithms using qubitization [29, 2, 44]. Its action is defined from Hamiltonian $H=\sum_{i=0}^{L-1}\alpha_{i}P_{i}$ as follows:
@@
In the following evaluation, we compare four methods: BFS, look-ahead BFS, Dijkstra projection, and look-ahead Dijkstra projection. We omitted the results of the 3D BFS and the 3D Dijkstra since they cannot finish the compilation within 1 hour.
@@
The Dijkstra projection and its optimized variant show better throughput than the two baseline methods. Compared to the look-ahead BFS, the look-ahead Dijkstra projection shows improvement by a factor of $1.6$ for SELECT 0 and $3.8$ for SELECT 5. We observed higher throughput improvement for SELECT 5, which is an expected result considering that SELECT 5 has a structure like a parallel circuit and leaves much room for parallelization. We can see that the degradation by considering kink conditions is at most 5%, which is not negligible, but does not lose the advantage of the Dijkstra projection.
@@
The order of execution times of the four methods for the SELECT circuits was basically the same as for the random circuit. On the other hand, the look-ahead Dijkstra projection demonstrated a significant speedup compared to the original Dijkstra projection when applied to SELECT 5. Further profiling indicated that these execution times were dominated by path searching, which strongly suggests that the instruction look-ahead technique results in substantially earlier termination of the path search subroutine. This is likely because fetching the instruction with the lowest height by instruction look-ahead flattens the current scheduling outcome with less biased height distribution, thereby shortening the subsequent lattice-surgery paths and the time required to find them.
All four methods finished in practical execution time. The additional time for considering kink conditions is at most 2%. Thus, the time for considering kink conditions is also negligible in this case. Also, when applied to the SELECT 0 circuit, the 3D BFS and the 3D Dijkstra cannot finish within an hour while the Dijkstra projection finishes within a few seconds, again indicating that the Dijkstra projection excels in execution time among the methods that utilize the time direction.
@@ 4.3 Comparison with the existing methods
In the following evaluation, four of our methods—BFS, look-ahead BFS, Dijkstra projection, and look-ahead Dijkstra projection—were compared with the Edge-Disjoint Paths Compilation (EDPC) algorithm [4]. We adopted a problem setting that satisfies the requirements of the EDPC algorithm. Specifically, the instruction sequence consists solely of CNOT gates and $T$-gates. Each $T$-gate is realized via $T$-gate teleportation between the target cell and a magic state factory cell aligned at the perimeter of the qubit plane. (For further details, see Appendix B.4.) Under this setting, SELECT 0 required $17$ qubit plane size and 37,85637{,}856 instructions, and SELECT 5 required $30$ qubit plane size and 39,05039{,}050 instructions.
Table 5: Throughput evaluation on SELECT circuits. The grayed rows annotated with an asterisk represent reference values that disregard the kink condition, optimistically assuming that CNOT operations complete via a 2D path in one code beat. Although these raw infeasible values are used to compare net routing performance, halving these grayed values will provide the corresponding feasible values, reflecting that CNOT operations realistically require two code beats. The remaining methods adhere to the kink condition and yield a feasible solution.
@@
The throughput of these methods is shown in Tab. 5. The grayed results represent configurations that violate the kink condition, rendering them infeasible. Since a CNOT gate can be implemented via a 2D path using two code beats [4], these entries can be converted to feasible counterparts by halving the values. However, the subsequent discussion compares the raw values to evaluate routing efficiency without this overhead. Notably, since our proposed methods satisfy the kink condition with modest overheads, accounting for the feasibility would effectively double the performance advantage, making our results even more favorable compared to the baseline methods.
For the performance of the EDPC algorithm, it lies between that of the BFS and the look-ahead BFS. This is due to their different strategy of managing instruction dependency. Their algorithm initially applies topological sorting to the instruction dependency graph and repeatedly removes the instructions that have no dependent instructions, thereby decomposing the instructions into a sequence of instruction sets consisting of mutually independent instructions. This partitioning enables a parallel compilation of the instruction sets and facilitates designing a scheduler that handles many types of instructions, whereas it also worsens the scheduling performance by effectively adding “barriers” in the circuit. We consider this to be one of the reasons why the look-ahead BFS outperforms EDPC, although their routing strategy is more sophisticated than our baseline methods.
Although the overall trends of our methods are consistent with the results in Sec. 4.2, it is worth mentioning that the instruction look-ahead technique proved exceptionally effective in this instance. This almost closed the gap between the look-ahead BFS and the Dijkstra projection, and resulted in an outstanding performance by the look-ahead Dijkstra projection.
This gap reduction can also be attributed to the difference in CNOT gate handling. The former experiment decomposes CNOT gates into $ZZ$ and $XX$ measurements, which partially yields stair-like structures and is likely advantageous for 3D routing. However, we emphasize that the same benefits can also be observed in this experiment when we compare feasible throughputs by halving the grayed ones: the Dijkstra projection showed $1.8$-fold improvement over the look-ahead BFS, and the look-ahead Dijkstra projection demonstrated $3.6$-fold improvement.
@@
The execution times of the methods are shown in Tab. 6. The substantial execution time of EDPC may stem from several reasons. We first note that their algorithm was implemented in Julia, whereas ours was implemented in C++. Although we used @btime to warm up the program and exclude just-in-time compilation time, this difference likely introduced overhead to their execution time. Another factor is their use of maximum flow subroutines in $T$-gate teleportation, which may have incurred high computational cost.